OpenStudy (anonymous):

the angles of a triangle are in the ratio 2:3:5. if the biggest side is 35cm

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

the biggest is 35, therefore the smallest is 14, and the middle is 21

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

find the measures of all the angles?

OpenStudy (anonymous):

 for all angles use law of cosines \[c^2=a^2+b^2-2ab \cos C\] \[a^2=c^2+b^2-2cb \cos A\] \[b^2=a^2+c^2-2ac \cos B\] Solve for the angles

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Can you get cos C by itself in the first equation?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Then you found C

OpenStudy (anonymous):

C=35?

OpenStudy (anonymous):

no, side c = 35 and C is different.

OpenStudy (anonymous):

the angles of a triangle are in the ratio 2:3:5. if the biggest side is 35cm

OpenStudy (anonymous):

 35cm isn't an angle

OpenStudy (anonymous):

a=21 b=14?

OpenStudy (anonymous):

those are the side lengths

OpenStudy (anonymous):

so what is the measures of all angles?

OpenStudy (anonymous):

how can i get the measures of the all angles?

OpenStudy (anonymous):

The Law of Cosines, like I wrote above \[c^2=a^2+b^2-2ab \cos C\] \[c^2-a^2-b^2=-2ab \cos C\] \[ \cos C = \frac{c^2-a^2-b^2}{-2ab}\] \[C = \arccos \Bigg(\frac{c^2-a^2-b^2}{-2ab} \Bigg) \]

OpenStudy (anonymous):

you do that for 2 of the angles, and can figure out the third through subtracting from 180º

OpenStudy (anonymous):

arccos?

OpenStudy (anonymous):

or inverse cosine on calculators it's sometimes written as \[cos^{-1}\]

OpenStudy (anonymous):

ok i solve it

OpenStudy (anonymous):

C=180